Determine the frequency , wavelength, speed and period in the equation y=2.5π cos(110t+ x/0.025)

1
Expert's answer
2022-01-05T11:37:26-0500

(a)

"f=\dfrac{\omega}{2\pi}=\dfrac{110\ \dfrac{rad}{s}}{2\pi}=17.5\ Hz."

(b)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{\dfrac{1}{0.025}\ m^{-1}}=0.157\ m."

(c)

"v=f\lambda=17.5\ Hz\times0.157\ m=2.75\ \dfrac{m}{s}."

(d)

"T=\dfrac{1}{f}=\dfrac{1}{17.5\ Hz}=0.057\ s."